OpenStudy (anonymous):

What is the simplified form of the SQUARE ROOT of 140

OpenStudy (anonymous):

2 root 35

OpenStudy (radar):

\[\sqrt{140}=\sqrt{4}\sqrt{35}=2\sqrt{35}\]